# A tale of two localizations: coexistence of flat bands and Anderson localization in a photonics-inspired amorphous system

### v1

@@ Abstract
Emerging experimental platforms use amorphousness, a constrained form of disorder, to tailor meta-material properties. We study localization under this type of disorder in a class of $2D$ models generalizing recent experiments on photonic systems. We explore two kinds of localization that emerge in these models: Anderson localization by disorder, and the existence of compact, macroscopically degenerate localized states as in many crystalline flat bands. We find localization properties to depend on the symmetry class within a family of amorphized kagomé tight-binding models, set by a tunable synthetic magnetic field. The flat-band-like degeneracy innate to kagomé lattices survives under amorphousness without on-site disorder. This phenomenon arises from the cooperation between the structure of the compact localized states and the geometry of the amorphous graph. For particular values of the field, such states emerge in the amorphous system that were not present on the kagomé lattice in the same field. For generic states, the standard paradigm of Anderson localization is found to apply as expected for systems with particle-hole symmetry (class D), while a similar interpretation does not extend to our results in the general unitary case (class A). The structure of amorphous graphs, which arise in current photonics experiments, allows exact statements about flat-band-like states, including such states that only exist in amorphous systems, and demonstrates how the qualitative behavior of a disordered system can be tuned at fixed graph topology.
@@ I Introduction
Rigorous results for amorphous systems are scarce, with the notable exception of the demonstrated persistence of a spectral gap in amorphous tetrahedral bonded semiconductors despite disorder [7, 8]. The robustness of the gap to amorphous disorder is markedly different from the case of random on-site disorder, where sharp crystalline bands broaden with increasingly strong disorder until the material eventually becomes a featureless insulator. In our work, we demonstrate and prove new exact results about the persistence of macroscopically degenerate states coming from compactly localized states in the presence of amorphous disorder. In a slight misappropriation of language, we refer to these persistent highly degenerate energy levels as “flat bands” even though there is no Brillouin zone; they are flat bands as a function of external fluxes through a torus [9]. Furthermore, quite remarkably, we show that flat bands exist in amorphous systems that have no analog in their crystalline counterparts. Flat bands are conducive to the formation of correlated states, such as superconductivity [10], quantum spin liquids [11] and the fractional quantum hall effect [12], and our model offers an interesting playground for this exploration as well as novel thermodynamic and transport features that accompany a spike in the density of states.
@@
There are two standard pathways to the formation of flat bands: compact localized states due to destructive interference of hopping paths on a geometrically frustrated lattice, such as the kagomé [13] or pyrochlore lattice [14], and the Landau-level-like mechanism leading to topologically nontrivial flat bands[15]. Both types of flat bands are generically destroyed by on-site and other traditionally studied forms of disorder. In the context of this work, we only encounter flat bands of the first type, which arise from compactly localized states.
@@
The mentioned decreased electron mobility in non-crystalline materials comes from the localization of some eigenmodes in the spectrum and it has also been observed for our model [5]. This phenomena is known as Anderson localization [16]. In the case of on-site disorder, a comprehensive paradigm has been established through application of thoroughly developed analytic tools [17] and tests in numerical simulations [18]. In amorphous systems the fate of Anderson localization has been considered [19, 20, 21, 22, 23] but a systematic study has not yet been undertaken. We begin to address this issue in this work by considering a model of amorphousness relevant to experiments in photonics. In particular, we find that, in some symmetry classes, for our model even the qualitative features of localization seem distinct from the standard paradigm.
@@
The rest of the work is organized as follows. In Sec. II we introduce the model, defined on a kagomé lattice, and explain how we construct amorphous structures that are locally identical to kagomé lattices. In Sec. III we explain why flat bands exist in some members of this family of model Hamiltonians and show how not only do these flat bands persist in the presence of amorphous disorder, there are in fact model Hamiltonians on amorphous structures that host flat bands which are absent in their periodic counterparts. We then turn to examining localization in the dispersive bands of our model. In Sec. IV we introduce a method to compute the effective localization length of eigenstates and adapt the method of energy level spacing distribution analysis to our amorphous systems. In Sec. IV.1 and IV.2 we compare the localization properties of model systems tuned to be in two different symmetry classes.
We discuss differences and challenges that this comparison raises to the conventional symmetry classification of localization in disordered systems.
@@ II.1 Amorphous Model
Beginning with an uncorrelated set of $L_{0}^{2}$ points in an $L_{0}\times L_{0}$ region we use kagomization [25] to generate a graph with the short range order defined above but no long range order. This process can be consistently done for any system size $L_{0}$ which is necessary for extrapolating the results to the thermodynamic limit, see as in Sec. IV. The kagomization procedure is explained in Fig. 1: we first construct a Voronoi diagram from the uncorrelated set of points, and then generate triangles around each Voronoi vertex by connecting the midpoints of all Voronoi edges emanating from the vertex. These triangles form a kagomized graph with $S\sim 6L_{0}^{2}$ sites. We define the system size as $L=\sqrt{S}$. To implement this graph with periodic boundary conditions, we perform the same procedure for a random point set on a two-dimensional torus.
@@
To write the model of Eq. (1) on a amorphous graph, we must be careful to only define hoppings around triangles that belonged to the original kagomé lattice. We refer to triangles of an amorphous graph that belonged to the original kagomé lattice as primitive triangles, and any cycles on the graph formed by connected primitive triangles as secondary polygons. On the kagomé lattice, all secondary polygons are hexagons, whereas on amorphous kagomized graphs, secondary polygons consist of $N$-sided polygons where $N=3,4,5,6,\cdots$. Writing our model on secondary triangles will give incorrect hopping terms. Thus, we define Eq. 1 on an amorphous graph by summing over all primitive triangles $\Delta_{i}$ of the graph. Numerically, this distinction is accomplished by writing the Hamiltonian from hoppings around the Voronoi vertices that give rise to a given kagomized graph.
@@ III.1 Robustness of flat bands to amorphousness in absence of field
In the kagomé lattice at $\phi$ = 0, the local value (non-normalized) of a given flat band state |$\ket{\psi_{fl}}$ alternates as ±$\pm 1$ around a single hexagon and vanishes elsewhere, see Fig. 2 (a). When the Hamiltonian acts on |$\ket{\psi_{fl}}$, destructive interference on each site neighboring the hexagon prevents this state from dispersing outside of the hexagon. This flat band state has degeneracy $S/3+1$.
@@
    The state has (un-normalized) amplitudes ±$\pm 1$ which alternate along the loop.
An example of a Type-I flat band state localized around a pentagon-septagon pair is shown in Fig. 2(b). Note that the wavefunction on the central site of this loop vanishes and that the action of the Hamiltonian on this state leads to destructive interference on this central site and all sites immediately neighboring the combined polygon. Type-I states can also reside on non-contractible loops of the torus.
@@
    The state has (un-normalized) amplitudes ±$\pm 1$ which alternate around the loops, but necessarily leaves two adjacent sites with the same sign, giving each loop a net sign
@@
    The connecting path must start at the third vertex of the primitive triangle that contains two adjacent sites with the same sign on one polygon, and end at the analogous vertex of the other polygon. The state on this path has (un-normalized) alternating amplitudes ±$\pm 2$
@@ III.2 Robustness of flat bands to amorphousness at non-zero field
Remarkably, for $p=1,2,6$ which correspond to $\phi=\pi/2,0,\pi/3$, we see that the fraction of all eigenstates that are in the flat band remains $1/3$ even on the amorphous graphs. In general we do not expect this to be the case. For example in Table 1 we see that for $\phi=\pi/6$, the fraction of flat band states never exceeds $1/16$ in our simulations. Additionally, we observe that this fraction varies appreciably from realization to realization for most values of $\phi$, see Table 1. The reasoning for why this does not occur for $p=2$ was explained in the previous subsection; the case $\phi=\pi/2$ ($p=1$) can also be easily understood. The flat band state in this case does not alternate ±$\pm 1$ around a polygon but is in fact uniform. This simple structure of the flat band states ensures compatibility with all secondary polygons in the graph: every secondary $n$-gon in an amorphous system supports a uniform localized flat band state surrounding it, see Fig. 2d. We believe an analysis similar to the previous subsection can be performed in these cases except that one would need to consider directed graphs. We leave this for future work.
Figure 2: Figures on left side represent a flat band state at $\phi=0$; figures on right side represent a flat band state at $\phi=\pi/2$. (a) Flat band eigenstate alternates between $+1$ (blue circles) and $-1$ (red circles) around the hexagon and is zero-valued outside. (b) Flat band eigenstate of an amorphous kagomized system wraps around a pentagon-septagon combination such that the state’s sign can still alternate between circumferential sites. (c) At $\phi=\pi/2$ the flat band state has the same phase at every site surrounding a hexagon. (d) Since the state’s sign is constant around a polygon, the pictured flat band state localized around a pentagon in an amorphous kagomized system is an eigenstate for $\phi=\pi/2$.
Figure 3: Example of a flat band state at $\phi=0$ that combines two non-adjacent polygons of odd length. The flat band eigenstate alternates between $+1$ (blue circles) and $-1$ (red circles), but leaves a net sign around each pentagon. While the compact localized state around each pentagon individually could not be an eigenstate, a “Dirac string" of $+2$ (blue heart) and $-2$ (red heart) connects the pentagons resulting in a flat band eigenstate.
@@ III.3 Flat bands unique to amorphous systems
Figure 4: Example of a case where a flat band exists in the amorphous system but not in the kagomé lattice system. For $\phi=11\pi/10$, a flat band emerges in the amorphous system in the gap of the kagomé lattice. This flat band contains .086.086.086.086 of the total states of the system, unlike the $1/3$ typically observed in flat bands in these systems. We hypothesize that these flat band states are hosted on pentagons.
@@ III.4 Other lattices
Figure 5: Example of a localized state on an amorphous analog of the Lieb lattice. In the Lieb lattice, states can exist both at vertices of the graph as well as at midpoints of the graph edges, denoted by $X$ in the figure. Blue (red) $X$ denote non-normalized state amplitude $+1$($-1$) whereas white circles at the vertices and black $X$s denote no state amplitude.
@@ IV Anderson-like localization in amorphous systems
Where $S$ is the number of sites on the graph and $\psi$ is normalized. We define an effective localization length $\tilde{\xi}=\frac{1}{\sqrt{I_{2}}}$. While inequivalent to the true localization length $\xi$, best approximated using the smallest quasi-1d Lyapunov exponent [33], the effective localization length is well suited to describing localization in amorphous systems for systems sufficiently small to be efficiently diagonalized.
We also analyze nearest-level spacings distributions. The set of nearest level spacings ${\Delta E}$ for a system is defined as the absolute difference between all nearest neighbors in an ordered set of the system’s eigenenergies. To define a distribution of the spacings, spacings are first normalized by the mean spacing to define $s=\Delta E/<\Delta E>$. In the limit of infinite system size, nearest-level spacings distributions in Anderson model disordered systems have been shown to fit either a Poisson law, $P_{P}(s)=1/\delta e^{-(s/\delta)}$, or the Wigner surmise for the orthogonal class ($\beta=1$), $P_{O}(s)=\frac{\pi}{2\delta}se^{-(\pi/4s^{2})}$ or a hybrid function, [34]. These three cases correspond to systems with only localized states, only delocalized (metallic) states, and systems at the metal-insulator transition.
@@
An important difference from the Anderson model is that, for $\phi\neq 0$, systems break time reversal symmetry. Thus in our analysis we use the Wigner surmise for the unitary class (class A), $\beta=2$, $P_{W}(s)=\frac{32}{\pi^{2}}s^{2}e^{-4s^{2}/\pi}$ [36]. The Wigner surmise for class D coincides with that of class A for states at sufficiently high energy [37]. We define subsets of all eigenstates using cutoff $\tilde{\xi}/L<1/3$ to approximate which states are localized, shown in Fig. 6, 8. We examine the largest system size only. For each amorphous system of this size, we choose all energies within a range $R$. Then we select energies ${E_{i,R}}$ in this range and find the nearest level spacings ${\Delta E_{i,R}}$ and normalize this set by its mean. We then compute a histogram for the normalized spacings ${\Delta E_{i,R}}$. Using the same histogram bins, we repeat this procedure for all realizations, then find the average value and error bar for each bin. For each range $R$, we fit a Poisson distribution and a Wigner distribution and then compute the $p$-value from the Kolmogorov-Smirnoff statistic $p$ for each fit. A large value $p\sim 1$ indicates that the data is likely to have come from this type of distribution whereas $p\sim 0$ indicates that the data comes from a different distribution. Results of these analyses are shown in Fig. 7, 9.
@@ IV.1 Anderson-like localization in class D
The effective localization spectrum for a large class D amorphous system ($L_{0}$ = 100) is shown in Fig. 6. The density of states of the bulk band extends into the bandgap of the kagomé lattice but the amorphous disorder is not sufficiently strong to close the gap. Most states in the amorphousness-induced tail, with energies $E<\sqrt{3}$, have suppressed effective localization length. However, not all states with $\tilde{\xi}<<L$ are in the tails: even in large systems, some states deep in the metallic (delocalized) region of the bulk are relatively localized. In fact, effective localization length $\tilde{\xi}$ can vary considerably between states that are close in energy, forming an envelope structure similar to that observed in the kagomé lattice. The envelope structure complicates defining a mobility edge in these systems. To approximate a mobility edge, we define energy ranges such that the effective localization length probability distribution can be split into two distinct regions, the “localized" states, shown in blue in Fig. 6 and “delocalized" states shown in red, with minimal overlap. All eigenstates outside of these categories we categorize as “metal insulator transition" states (purple). Setting cutoff $\tilde{\xi}/L<1/3$ for classifying localized states coincides with this analysis.
Figure 6: Top: Scatter plot of the energy of every eigenstate in the highest energy band of an $L_{0}=100$ amorphous system and its effective localization length. States are separated into three energy ranges, corresponding to localized (blue), metal-insulator-transition (purple) and delocalized (red). Bottom: States classified as localized (red) have nonzero probability $P(\tilde{\xi}/L)$ for $\tilde{\xi}/L<1/3$. The energy range for delocalized states is set such that the delocalized states’ (red) probabilities’ non-zero range has minimal overlap with the localized states’ range. All states in-between are classified as metal insulator transition states, whose distribution $P(\tilde{\xi}/L)$ is shown in purple.
@@
The nearest-level-spacings of states assigned to the localized energy range fit a Poisson law well. The spacings of states assigned to the delocalized energy range do not come from a Poisson law distribution, nor do they fit the Wigner distribution with a high degree of certainty. This inconclusivity is likely due to the persistence of localized states deep in the metallic phase, as shown in Fig. 6.

### v2

@@ Abstract
Emerging experimental platforms use amorphousness, a constrained form of disorder, to tailor meta-material properties. We study localization under this type of disorder in a family of 2D models generalizing recent experiments on photonic systems. Models in this family reside on amorphous analogs of kagomé lattices with fixed coordination number, vary by a tunable synthetic field, and remarkabaly, permit exact results. We observe two kinds of localization that emerge in these models: Anderson localization by amorphous disorder, and the existence of compact, macroscopically degenerate localized states as in many crystalline flat bands. The flat-band-like degeneracy innate to kagomé lattices survives under amorphousness without on-site disorder. This phenomenon arises from the cooperation between the structure of the compact localized states and the geometry of the amorphous graph. More surprisingly, for particular values of the field, such states emerge in the amorphous system that were not present on the kagomé lattice in the same field. Outside the flat band, constrained amorphous graph geometry necessitates the existence of a fully delocalized state, near which we observe evidence of a localization-delocalization transition. Our platform serves as a demonstration of how the qualitative behavior of a disordered system can be tuned at fixed graph topology and lead to localization phenomena unique to amorphous systems that are not observed in their generically disordered counterparts.
@@ I Introduction
Rigorous results for amorphous systems are scarce, with the notable exception of the demonstrated persistence of a spectral gap in amorphous tetrahedral bonded semiconductors despite disorder [7, 8]. The robustness of the gap to amorphous disorder is markedly different from the case of random on-site disorder, where sharp crystalline bands broaden with increasingly strong disorder until the material eventually becomes a featureless insulator. In our work, we demonstrate and prove new exact results about the persistence of macroscopically degenerate states coming from compactly localized states in the presence of amorphous disorder. In a slight misappropriation of language, we refer to these persistent highly degenerate energy levels as “flat bands” even though there is no Brillouin zone; they are flat bands as a function of external fluxes through a torus [9]. For general reviews of results on flat bands, see [10, 11, 12, 13, 14, 15, 16]. Furthermore, quite remarkably, we show that flat bands exist in amorphous systems that have no analog in their crystalline counterparts. Flat bands are conducive to the formation of correlated states [17], such as ferromagentism [18], superconductivity [19], quantum spin liquids [20] and the fractional quantum hall effect [21], and our model offers an interesting domain for such exploration as well as novel thermodynamic and transport features that accompany a spike in the density of states.
@@
There are two standard pathways to the formation of flat bands [13]: compact localized states due to destructive interference of hopping paths on a geometrically frustrated lattice, such as the kagomé [22] or pyrochlore lattice [23], and the Landau-level-like mechanism leading to topologically nontrivial flat bands[24]. In the context of this work, we only encounter flat bands with compactly localized states (the first type). Flat bands are generically destroyed by on-site and other traditionally studied forms of disorder, although the states can stay localized [25]. Quite remarkably, it was noticed that disorder that preserves local connectivity [26, 27, 28] as well other types of specially engineered breaking of periodicity [29, 30, 31, 15] can leave flat bands and compact localized states therein intact. We observe similar phenomena for a general class of experimentally relevant amorphous models with partial or full preservation of the flat band degeneracy and non-trivial modification of the geometry of localized states.
@@
Compact localized states are not the only localized states we observe in these models. Localization also originates from destructive interference of eigenmodes across randomized scattering paths. This phenomenon, known as Anderson localization [32], has previously been observed for a class of our model [5]. In the case of on-site disorder, a comprehensive paradigm has been established through application of thoroughly developed analytic tools [33] and tests in numerical simulations [34]. In amorphous systems the fate of Anderson localization has been considered [35, 36, 37, 38, 39] but a systematic study has not yet been undertaken. We begin to address this issue in this work by considering a model of amorphousness relevant to experiments in photonics. In particular, we find that, in some symmetry classes, for our model even the qualitative features of localization seem distinct from the standard paradigm.
@@
The rest of the work is organized as follows. In Sec. II we introduce the model, defined on a kagomé lattice, and explain how we construct amorphous structures that are locally identical to kagomé lattices. In Sec. III we explain why flat bands exist in some members of this family of model Hamiltonians and show how not only do these flat bands persist in the presence of amorphous disorder, there are in fact model Hamiltonians on amorphous structures that host flat bands which are absent in their periodic counterparts. We then turn to examining localization in the dispersive bands of our model. In Sec. IV we introduce a method to compute the effective localization length of eigenstates and adapt the method of energy level spacing distribution analysis to our amorphous systems. In Sec. IV.1 and IV.2 we compare the localization properties of model systems tuned to be in two different symmetry classes. Finally, in Sec. IV.3, we discuss our observation of a localization-delocalization transition unique to amorphous systems.We explain how this behavior arises because of properties beyond the conventional symmetry classification of localization in disordered systems, chiefly the coexistence of randomness with constrained local coordination in amorphous systems.
@@ II.1 Amorphous Model
Beginning with an uncorrelated set of $L_{0}^{2}$ points in an $L_{0}\times L_{0}$ region we use the line graph construction [18] to generate a graph with the short range order defined above but no long range order. Following [41], we refer to this procedure as kagomization due to the local similarity between the resulting graphs and the kagomé lattice. This process can be consistently done for any system size $L_{0}$ which is necessary for extrapolating the results to the thermodynamic limit, see as in Sec. IV. The kagomization procedure is explained in Fig. 1: we first construct a Voronoi diagram from the uncorrelated set of points, and then generate triangles around each Voronoi vertex by connecting the midpoints of all Voronoi edges emanating from the vertex. These triangles form a kagomized graph with $S\sim 6L_{0}^{2}$ sites. We define the system size as $L=\sqrt{S}$. To implement this graph with periodic boundary conditions, we perform the same procedure for a random point set on a two-dimensional torus.
@@
To write the model of Eq. (1) on a amorphous graph, we must be careful to only define hoppings around triangles that belonged to the original kagomé lattice. We refer to triangles of an amorphous graph that belonged to the original kagomé lattice as primitive triangles, and any cycles on the graph formed by connected primitive triangles as secondary polygons. On the kagomé lattice, all secondary polygons are hexagons, whereas on amorphous kagomized graphs, secondary polygons consist of $N$-sided polygons where $N=3,4,5,6,\cdots$. Writing our model on secondary triangles will give incorrect hopping terms. Thus, we define Eq. 1 on an amorphous graph by summing over all primitive triangles $\Delta_{i}$ of the graph. Numerically, this distinction is accomplished by writing the Hamiltonian from hoppings around the Voronoi vertices that give rise to a given kagomized graph. For example, to define the model on the graph shown in Fig. 1, we assign hoppings to all of the black edges. The signs the hopping phases are determined by their positions in the triangles of black edges that encircle the orange points.
@@ III.1 Robustness of flat bands to amorphousness in absence of field
[…]
@@
    The state has (un-normalized) amplitudes $\pm 1$ which alternate along the loop.
An example of a Type-I flat band state localized around a pentagon-septagon pair is shown in Fig. 2(b). Note that the wavefunction on the central site of this loop vanishes and that the action of the Hamiltonian on this state leads to destructive interference on this central site and all sites immediately neighboring the combined polygon. Type-I states can also reside on non-contractible loops of the torus. For periodic systems, the presence of non-contractible loops among flat band states was previously recognized in [13].
@@
    The state has (un-normalized) amplitudes $\pm 1$ which alternate around the loops, but necessarily leaves two adjacent sites with the same sign, giving each loop a net sign
@@
    The connecting path must start at the third vertex of the primitive triangle that contains two adjacent sites with the same sign on one polygon, and end at the analogous vertex of the other polygon. The state on this path has (un-normalized) alternating amplitudes $\pm 2$
@@ III.2 Robustness of flat bands to amorphousness at non-zero field
Remarkably, for $p=1,2,6$ which correspond to $\phi=\pi/2,0,\pi/3$, we see that the fraction of all eigenstates that are in the flat band remains $1/3$ even on the amorphous graphs. In general we do not expect this to be the case. For example in Table 1 we see that for $\phi=\pi/6$, the fraction of flat band states never exceeds $1/16$ in our simulations. Additionally, we observe that this fraction varies appreciably from realization to realization for most values of $\phi$, see Table 1. The reasoning for why this does not occur for $p=2$ was explained in the previous subsection; the case $\phi=\pi/2$ ($p=1$) can also be easily understood. The flat band state in this case does not alternate $\pm 1$ around a polygon but is in fact uniform. This simple structure of the flat band states ensures compatibility with all secondary polygons in the graph: every secondary $n$-gon in an amorphous system supports a uniform localized flat band state surrounding it, see Fig. 2d. We believe an analysis similar to the previous subsection can be performed in these cases except that one would need to consider directed graphs. We leave this for future work.
Figure 2: Figures on left side represent a flat band state at $\phi=0$; figures on right side represent a flat band state at $\phi=\pi/2$. (a) Flat band eigenstate alternates between +1+1 (blue circles) and 1-1 (red circles) around the hexagon and is zero-valued outside. (b) Flat band eigenstate of an amorphous kagomized system wraps around a pentagon-septagon combination such that the state’s sign can still alternate between circumferential sites. (c) At $\phi=\pi/2$ the flat band state has the same phase at every site surrounding a hexagon. (d) Since the state’s sign is constant around a polygon, the pictured flat band state localized around a pentagon in an amorphous kagomized system is an eigenstate for $\phi=\pi/2$.
Figure 3: Example of a flat band state at $\phi=0$ that combines two non-adjacent polygons of odd length. The flat band eigenstate alternates between +1+1 (blue circles) and 1-1 (red circles), but leaves a net sign around each pentagon. While the compact localized state around each pentagon individually could not be an eigenstate, a “Dirac string" of +2+2 (blue heart) and 2-2 (red heart) connects the pentagons resulting in a flat band eigenstate.
@@ III.3 Flat bands unique to amorphous systems
Figure 4: Example of a case where a flat band exists in the amorphous system but not in the kagomé lattice system. For $\phi=11\pi/10$, a flat band emerges in the amorphous system in the gap of the kagomé lattice. This flat band contains .086.086 of the total states of the system, unlike the $1/3$ typically observed in flat bands in these systems. We hypothesize that these flat band states are hosted on pentagons.
@@ III.4 Other lattices
Figure 5: Example of a localized state on an amorphous analog of the Lieb lattice. In the Lieb lattice, states can exist both at vertices of the graph as well as at midpoints of the graph edges, denoted by $X$ in the figure. Blue (red) $X$ denote non-normalized state amplitude +1+1(1-1) whereas white circles at the vertices and black $X$s denote no state amplitude.
@@ IV Anderson-like localization in amorphous systems
Where $S$ is the number of sites on the graph, $\psi_{i}$ is the amplitude of the state on site $i$, and $\psi$ is normalized. We define an effective localization length $\tilde{\xi}=\frac{1}{\sqrt{I_{2}}}$. While inequivalent to the true localization length $\xi$, best approximated using the smallest quasi-1d Lyapunov exponent [48], the effective localization length is well suited to describing localization in amorphous systems for systems sufficiently small to be efficiently diagonalized.
We also analyze nearest-level spacings distributions, adapted for these systems. In the main text we give the results, specifically whether the distribution fits a Poisson law (localized states), a Wigner surmise (delocalized states), or neither. Details of these analyses are described in App. B.
@@ IV.1 Anderson-like localization in class D
The effective localization spectrum for a large class D amorphous system ($L_{0}$ = 100) is shown in Fig. 6. The density of states of the bulk band extends into the bandgap of the kagomé lattice but the amorphous disorder is not sufficiently strong to close the gap. Most states in the amorphousness-induced tail, with energies $E<\sqrt{3}$, have suppressed effective localization length. However, not all states with $\tilde{\xi}<<L$ are in the tails: even in large systems, some states deep in the metallic (delocalized) region of the bulk are relatively localized. In fact, effective localization length $\tilde{\xi}$ can vary considerably between states that are close in energy, forming an envelope structure similar to that observed in the kagomé lattice. The envelope structure complicates defining a mobility edge in these systems. To approximate a mobility edge, we define energy ranges such that the effective localization length probability distribution can be split into two distinct regions, the “localized" states, shown in blue in Fig. 6 and “delocalized" states shown in red, with minimal overlap. All eigenstates outside of these categories we categorize as “intermediate" states (purple). Setting cutoff $\tilde{\xi}/L<1/3$ for classifying localized states coincides with this analysis.
Figure 6: Top: Scatter plot of the energy of every eigenstate in the highest energy band of an $L_{0}=100$ amorphous system and its effective localization length. States are separated into three energy ranges, corresponding to localized (blue), intermediate (purple) and delocalized (red). Bottom: States classified as localized (red) have nonzero probability $P(\tilde{\xi}/L)$ for $\tilde{\xi}/L<1/3$. The energy range for delocalized states is set such that the delocalized states’ (red) probabilities’ non-zero range has minimal overlap with the localized states’ range. All states in-between are classified as intermediate states, whose distribution $P(\tilde{\xi}/L)$ is shown in purple.
Nearest-level-spacings analysis of states assigned to the localized energy range fit a Poisson law well. The spacings of states assigned to the delocalized energy range do not come from a Poisson law distribution, nor do they fit the Wigner distribution with a high degree of certainty. These distributions are shown in Fig. 11 in App. B. This inconclusivity is likely due to the persistence of localized states deep in the metallic phase, as shown in Fig. 6.